### v2

# What Lies Beneath? Exploring the Impact of Underlying AI Model Updates in AI-Infused Systems

Abstract.
AI models are constantly evolving, with new versions released frequently. This raises a key question: how should AI-infused systems integrate updates when the downstream impact on user experience and performance is unclear? Human-AI interaction guidelines encourage notifying users about (changes in) model capabilities, ideally supported by thorough benchmarking. Yet, as AI models integrate into domain-specific workflows, exhaustive benchmarking can become impractical or expensive, often leading to invisible or minimally communicated updates. In this work, we explore the impact of such updates through two complementary studies on facial recognition for historical person identification. First, we conducted an online experiment to understand how users distinguish between models, followed by a diary study examining user perceptions in a real-world deployment. Our findings reveal how model changes impact human-AI performance, downstream user behavior, and the folk theories they develop. Based on these insights, we discuss implications for updating models in AI-infused systems.
AI model update, facial recognition, user behavior, user perception of AI models, information retrieval, transparency, human-AI interaction guidelines, historical person identification
1. Introduction
In September 2023, Tesla rolled out a new beta version of its Full Self-Driving (FSD) software, promising advanced autonomous capabilities. Yet, for one driver, the update took an unexpected turn. While cruising on the highway, the car abruptly swerved toward a median at high speed—forcing the driver to intervene to avoid a crash. Unfortunately, it was not an isolated case; several other drivers reported similar issues, leading to widespread confusion about what had changed. (Cao, 2023)
AI model developers are pushing out new models and updates at a rapid pace, driven by growing user needs and the necessity to incorporate newer capabilities and address inaccuracies. For instance, Hugging Face hosts over 300,000 models (Tousignant, 2023), and in 2023 alone, 149 new foundation models were released (Maslej et al., 2024). These models power a wide range of AI-infused applications across domains such as healthcare, finance, social media, and autonomous vehicles (Maslej et al., 2024). As these underlying models get updated, a critical question arises: how should the applications that build upon them incorporate these changes, and how do these updates impact the downstream user experience?
Previous work has shown that model updates do not always improve human-AI team performance (Bansal et al., 2019b; Chattopadhyay et al., 2017) and can lead to user dissatisfaction due to misaligned expectations or disruptions in user workflows (Kocielnik et al., 2019; Bucher, 2012). Such unintended consequences can manifest in various ways, be it unforeseen accidents following updates to autonomous vehicle algorithms (Cao, 2023), or distress among social media users due to alterations in recommendation algorithms (Register et al., 2023), or diminished volunteer engagement in citizen science projects upon the integration of automated systems (Trouille et al., 2019). To avoid unexpected shifts in expectations, human-AI interaction guidelines advocate for notifying users about model changes and capabilities (Amershi et al., 2019).
Effectively communicating an AI model’s capabilities requires a deep understanding of what the model can do, which in turn demands comprehensive benchmarking and stress testing across diverse scenarios. However, as these models are deployed in workflows or domains that developers did not originally anticipate or design for, the complexity of thorough benchmarking increases (Raji et al., 2021). This makes comprehensive testing both impractical and costly, leading to silent updates or minimal communication or approximation about changes and capabilities, which leaves users unaware of significant model modifications. When users are unaware of these changes, they may interact with the system based on outdated assumptions, leading to potential errors, inefficiencies, or frustrations (Bansal et al., 2019a; Villareale et al., 2022).
With the ubiquity of AI-infused applications, many lay end-users may not fully understand AI model capabilities, let alone grasp what has changed following updates. This lack of awareness raises key questions about how well users can adapt to these changes and whether they can complete their tasks effectively. In this work, we aim to understand whether users of an AI-infused application can detect changes in the underlying models in such practical scenarios with limited communication, and how these updates affect their performance, perceptions, and behavior.
To explore these questions, we conducted two complementary studies involving a facial recognition-based system for completing historical person identification tasks (i.e., finding potential matches from a ranked list of search results for a given query photo): 1) a controlled experiment where users were unaware of the underlying face recognition model being updated, and 2) a real-world deployment where they could toggle between the old and new models.
For the first study, we ran an online experiment to determine whether users could distinguish changes between facial recognition models across successive trials and what cues they relied on to do so. We also explored how the underlying model impacted their task performance, perceptions of accuracy, and behavior, particularly when they were unaware of the model change. We complemented this with a second study in a real-world setting on Civil War Photo Sleuth (CWPS) (Mohanty et al., 2019b), a widely-used platform for identifying historical photos through a facial recognition-based workflow. In April 2023, CWPS added a new facial recognition model alongside the existing one, allowing users to toggle between the two models. Through a two-week long diary study with 10 active users, we aimed to understand how they perceived the new model in contrast to the old one, and what factors influenced their model preferences.
Our findings provide key insights into users’ abilities to detect model changes, the cues they rely on, and how these changes influence their task performance, perceptions of accuracy, behavior, and the folk theories and preferences they develop. We discuss the complexities involved in end-user evaluation of AI models and the implications for deploying model updates in a human-AI collaborative environment.
2. Related Work
2.1. User Frustration with Software Updates
Software updates are ubiquitous in modern life, from operating system patches to app updates across devices. Most updates happen silently, with users often scheduling them for convenience, like overnight updates. However, some updates can have major, unforeseen consequences, such as the 2024 CrowdStrike incident that briefly disrupted global operations (Wikipedia contributors, 2024), or stock exchange glitches caused by a system update (Targett, 2023). At the consumer level, prior work by Vaniea et al. (Vaniea and Rashidi, 2016) shows that users go through a six-stage process when updating software, and unexpected changes, such as interface or functionality shifts, often lead to negative perceptions. Users frequently avoid updates when they anticipate disruptions to their workflows or loss of preferred features.
Though experts recommend auto-updates for security, many users disable them after experiencing issues like performance drops, preferring control over updates (Mathur and Chetty, 2017). Haney et al. (Haney and Furman, 2023) found that smart home users frequently face confusion from unclear update notifications and compatibility issues, suggesting that clearer communication and more control, such as manual vs. automatic updates, would improve the experience. Mathur et al. (Mathur et al., 2018) further emphasize that users avoid updates due to perceived costs (e.g., time, storage, risks like data loss), with unclear messaging exacerbating this reluctance. Morreale et al. (Morreale and Eriksson, 2020) further illustrate how software updates can drastically alter user experiences, as seen with Spotify’s 2019 update, which removed features and disrupted user workflows. This update led to widespread frustration as users lost control over how they organized and accessed their music, highlighting how updates can impose new norms of use that reduce user agency.
This resistance to updates is reflected in the findings from Rula et al. (Rula et al., 2020), which showed that many users continue to run outdated software, avoiding updates due to concerns about potential disruptions or changes. This behavior further demonstrates how user apprehension can hinder the adoption of updates, leading to increased security vulnerabilities and reduced system performance. While these studies highlight software updates, they parallel our focus on AI model updates, which further add unpredictability, especially with generative AI, making it critical to understand downstream effects with varying levels of user communication.
2.2. User Perceptions of Dynamic AI Systems
Advances in AI research have opened new avenues for human-AI collaboration, leveraging the complementary strengths of humans and AI systems to drive progress across diverse domains such as creativity, manufacturing, and journalism (Wilson and Daugherty, 2018; Liu et al., 2023; Broussard et al., 2019). However, this collaboration is far from static or fully understood. Both human users and AI agents operate dynamically, with interactions shaped by factors such as context of use, task complexity, and the continually evolving—yet often uncertain—nature of AI technologies (Yang et al., 2020; Amershi et al., 2019; Horvitz, 1999). These complexities are further amplified by the growing push toward Agentic AI, where large language model (LLM) agents are increasingly designed to act autonomously and make independent decisions (Wu et al., 2023; Maslej et al., 2024).
While the fusion of human expertise with AI capabilities has led to significant advancements, it is crucial to recognize the inherent complexities and potential drawbacks that accompany this collaboration. Dynamic human-AI interactions can give rise to challenges such as mistrust, misinformation, and data privacy concerns (Zhou et al., 2023; Spitale et al., 2023). In AI-assisted decision-making (Lai et al., 2023), additional pitfalls include automation bias, amplification of incorrect results, and the perpetuation of existing biases, which can further hinder harmonious collaboration (Willever-Farr and Forte, 2014; Jones-Jang and Park, 2023; Kim et al., 2024).
Prior research strongly advocates for a nuanced understanding of both human and AI capabilities to foster more effective collaboration (Bansal et al., 2019a; Wang et al., 2021; Cai et al., 2019; He and Gadiraju, 2022; Liao and Varshney, 2021). A substantial body of literature has explored the factors shaping user perception of AI, including algorithm speed (Park et al., 2019; Teevan et al., 2013; Gnewuch et al., 2018), the number of results generated (Oulasvirta et al., 2009), model explanations (Kim et al., 2023; Wang and Yin, 2021, 2023; Vasconcelos et al., 2023), trust calibration interventions (Ma et al., 2023; Buçinca et al., 2021), group behavior (Chiang et al., 2023), and the system’s ability to communicate its expertise effectively (Zhang et al., 2022). Research also shows that providing users with some degree of control can enhance their perception of and trust in the system (Kocielnik et al., 2019).
As the market experiences a rapid influx of new AI models, this frequent turnover, while often beneficial, presents unique challenges. One of the key questions is how users perceive and adapt to changes in AI systems, particularly when updates to the underlying models are not clearly communicated. Prior work has explored how users may perceive shifts in model explanations after updates, even when the decisions remain unchanged (Wang and Yin, 2023). These shifts can affect users’ trust, particularly if the new explanations diverge significantly from previous ones. Users may experience cognitive dissonance when their established mental models are violated, leading to reduced trust and satisfaction with the system.
However, there remains a notable gap in fully understanding user perceptions, particularly concerning updates in black-box models deployed in real-world scenarios. Prior research indicates that the introduction of new models is not always welcomed by users. In some cases, it can lead to decreased engagement and increased distress due to the unpredictability of the system’s behavior (Register et al., 2023; Trouille et al., 2019). This highlights the need for a more comprehensive understanding of how users react to major model upgrades, especially in systems where the internal workings are opaque and explanations cannot be easily communicated.
In this work, we explore how users perceive AI model updates both in a controlled environment and a real-world deployment, focusing on factors such as response speed, the number of results generated, and the cognitive load required to interpret changes. By examining these elements under different levels of communication about model updates, we aim to provide practical validation for existing human-AI interaction guidelines (Amershi et al., 2019),
2.3. Civil War Photo Sleuth and Historical Person Identification
Civil War Photo Sleuth (CWPS) [^1] https://www.civilwarphotosleuth.com/ is an online platform designed to help users identify unknown individuals in historical photos through a facial recognition-based pipeline (Mohanty et al., 2019b). By presenting a list of visually similar candidates, CWPS supports users in performing side-by-side facial similarity comparisons, a fairly generic visual inspection task (Mohanty and Luther, 2023). The platform’s established user base is already accustomed to interacting with AI-driven face recognition, making it an ideal setting for studying user perceptions of AI model updates. In the context of the historical person identification task, users rely on both the AI’s recommendations (i.e., search pool of facially similar-looking candidates) and their own subjective judgment to make decisions (i.e., find potential matches from the search pool), offering a unique opportunity to observe how different models influence decision-making. With over 50,000 photos in its database, CWPS provides a rich platform for investigating how users respond to varying model outputs.
For the first study, we developed a web platform that closely mirrored CWPS’s search interface and used their API to retrieve facially similar candidates for running experiments on Prolific. In the second study, we recruited 10 active CWPS users to compare the outputs of two distinct facial recognition models. We will explore the design, methodology, and findings of these studies in detail in the following sections.
3. Study 1: Distinguishing AI model updates without explicit communication
In this study, we investigate how updates in underlying AI models influence user perceptions, performance, and behaviors in environments where changes occur without explicit communication. To explore this, we conducted an online experiment where participants completed multiple trials of a historical person identification task. In each trial, they were asked to find matches for a given query photo by interacting with the results of a randomly assigned black-box facial recognition model.
Participants were not informed about which model they were using or if the model had changed between trials. Through this study, we examine 1) whether users can accurately detect changes in the underlying model, 2) what kind of cues they rely on to make this distinction, and 3) how their perceptions, performance, and behaviors shift when the model changes.
3.1. Hypotheses
Based on prior work and the distinctive differences between the two models (see Section 3.4), we propose the following hypotheses to guide our investigation:
  1. H1
    Users will be able to distinguish a change in the model even though they are interacting with black-box systems.
    Prior work has shown that users can detect subtle shifts in system behavior, though this has primarily been tested with model explanations in non-black-box, controlled settings (Wang and Yin, 2023).
  2. H2
    Users will rely on observable model characteristics such as the number of search results and latency to distinguish between models.
    Research shows that users often focus on tangible factors like speed and the volume of results while interacting with AI models (Park et al., 2019; Oulasvirta et al., 2009).
  3. H3
    Users will perceive the more accurate model (newer) as more accurate, even if they are not explicitly informed about model changes.
    Drawing from different cognitive science theories, users are likely to perceive the model that retrieves more similar-looking candidates as more accurate because they can detect subtle improvements in output quality by focusing on high-diagnostic and directly comparable features (Markman and Gentner, 1996; Tversky, 1977).
  4. H4
    Human-AI team performance will improve with the more accurate model.
    Higher model accuracy contributes to reduced errors, which in turn improves overall team outcomes (Towler et al., 2023). However, prior work has also shown that may not always be true (Bansal et al., 2019b; Chattopadhyay et al., 2017).
  5. H5
    The underlying model will affect user behavior in terms of how many results they check and how they make decisions.
    Changes in recommender algorithms, as seen in social media platforms, have been shown to significantly alter user behavior, leading them to engage more or less with certain types of content (Owen, 2019; Karizat et al., 2021). Users adjust their interaction patterns based on perceived changes in output quality, relevance, or volume, influencing their decision-making and exploration of results.
  6. H6
    When users believe that a model has switched, it will lead to changes in their behavior.
    According to Cognitive Dissonance Theory (Festinger, 1957), when users experience a mismatch between their expectations and system behavior, they adjust their actions to reduce the discomfort caused by this inconsistency, leading to changes in their decision-making and interaction strategies.
3.2. Experiment Setup
To address these questions, we designed an online survey experiment where participants interacted with a facial recognition-based web application (see Figure 1), modeled after the search results page on CWPS (Mohanty et al., 2019b). In each trial, participants were shown a query photo and tasked with identifying potential matches (i.e., other photos of the same person) from a ranked list of search results retrieved by a black-box facial recognition model. Participants visually inspected the results to determine whether any of the retrieved images matched the person in the query photo.
Each participant completed 8 trials, with each trial featuring a different query photo and its corresponding search results (i.e., potential candidates). We randomized the underlying model used in each trial (i.e., the model could either remain the same or change between trials). Participants were not informed of any model changes, allowing us to investigate whether they could detect changes and what cues they relied on. This design enabled us to examine shifts in user perceptions, performance, and behaviors across trials.
3.3. Experiment Interface
We designed the web application to guide participants through the person identification tasks across eight consecutive trials. The interface supported five key components:
  • Exploring Search Results: In each trial, participants were presented with a ranked list of potential candidates retrieved by a facial recognition model for a given query photo. The list was ordered by the model’s confidence score, which indicated how likely each candidate resembled the query photo (see Figure 1). Participants visually inspected this list of candidates to find possible matches for the query photo.
    Figure 1. Search interface for examining potential matches retrieved by the facial recognition model. The search results (right) are ranked according to the model’s confidence score, which ranges from 0% (not at all similar-looking) to 100% (highly similar-looking). Only candidates with confidence scores above 50% were retrieved, following the CWPS threshold. Each search result includes a unique photo ID (from CWPS), thumbnail of the face, and the model’s confidence score, quantifying its similarity to the query photo (on the left). Participants were required to compare at least three search results before completing the analysis.
[…]
    A search interface displaying a historical query photo on the left and a list of potential matches on the right, each accompanied by a confidence score. The list is ranked by confidence scores, ranging from 54.78% to 59.25%, indicating how similar each candidate is to the query photo. The interface requires participants to compare at least three results before completing the analysis.
  • Comparing Results: Participants can perform a close inspection on any search result by opening an interface that allows them to do a side-by-side comparison of the query photo and the selected search result (similar to (Mohanty and Luther, 2023)). This comparison interface allows users to then vote on whether the two photos showed a facial match (i.e., same person, different view), a replica (i.e., same person, same view), different people, or if they were not sure (see Figure 2).
    Figure 2. Side-by-side comparison interface for close inspection of a search result (right) with the query photo (left). Participants can carefully compare the two images and vote on one of four options: facial match (same person, different view), replica (same person, same view), different people, or not sure.
[…]
    Side-by-side comparison interface showing a historical query photo on the left and a search result on the right. Participants can choose from four options after comparing the two photos: facial match, replica, different people, or not sure.
  • Rating Model Accuracy: At the end of each trial (i.e., after participants finished comparing the search results), they rated the accuracy of the facial recognition model on a slider ranging from -100 (very inaccurate) to 100 (very accurate), with an optional text box to explain their rating (see Figure 3). We chose this wide range to give participants more flexibility to express shifts in perceived accuracy across trials, without attributing precise or fixed meaning to any specific point on the scale.
    Figure 3. Model rating interface where participants rated the accuracy of the facial recognition model used in the current trial. The slider ranged from -100 (extremely inaccurate) to 100 (extremely accurate). An optional text box was provided for participants to explain their rating, allowing them to elaborate on the factors that influenced their decision.
[…]
    A model rating interface showing a slider for participants to rate the accuracy of the facial recognition model. The scale ranges from -100 (extremely inaccurate) to 100 (extremely accurate), with an optional text box below to provide additional comments on the rating.
  • Comparing Models Across Trials: After rating the current model, participants compared the model used in the current trial with the one from the previous trial (see Figure 4). The interface displayed relevant trial-specific details such as the query photo, model latency (i.e., time taken to retrieve results), the number of search results, whether the participant found a match, and their perceived accuracy. Participants then rated the similarity or difference between the models on a four-point Likert scale: very different, somewhat different, somewhat similar, or very similar. The system automatically labeled the model used in the first trial as Model A. Based on the participants’ comparison responses, the system assigned models in the subsequent trials as either Model A or Model B.
    Figure 4. Interface for comparing the model used in the current trial with the model from the previous trial. The interface displays trial-specific details, including the query photo, model latency, number of search results, whether the participant found a match, and the participant’s perceived accuracy. Participants then rate the similarity or difference between the models on a four-point Likert scale: very different, somewhat different, somewhat similar, or very similar.
[…]
    Interface for comparing the models used in two consecutive trials. It shows the query photos, task summaries, model latency, number of search results, and perceived accuracy for each trial. Participants are asked to rate how similar or different the models used in the trials are.
  • Selecting a Preferred Model: After all 8 trials were completed, participants were shown a summary of their trial categorizations (see Figure 5), displaying which trials were categorized as Model A or Model B. Participants then selected their final preferred model for future photo identification tasks and provided a justification for their choice.
    Figure 5. Summary of all trials, categorized into Model A and Model B as perceived by the user. The first trial was automatically assigned to Model A, while the remaining trials were categorized based on the participant’s model comparison responses. The interface displays trial-specific information such as task duration, number of search results, and the participant’s accuracy rating. After viewing the summary, participants were asked to select their preferred model for future photo identification tasks and provide justification for their choice.
[…]
    A summary interface showing all trials categorized into Model A and Model B based on user perception. The first trial is assigned to Model A by default, while the remaining trials are categorized according to the participant’s comparison responses. Each trial displays details such as task duration, number of search results, and the participant’s accuracy rating.
In order to distinguish them from the actual underlying models (which were unknown to the participants), we will refer to these as Model A-UP and Model B-UP (user-perceived).
3.4. Facial Recognition Models
Similar to CWPS (Mohanty et al., 2019b), we used facial recognition models from Microsoft Azure for this study: the oldest available model, recognition_01 (released in 2017), and the most recent model, recognition_04 (released in 2021). We will refer to these as the old model and new model, respectively. According to the developer website (Urban, 2023), the new model is supposed to be the most accurate available, designed to handle challenging cases like faces with masks or facial hair. However, its effectiveness in specialized domains, such as historical photo identification, remains unclear. Previous benchmarking of the old model on CWPS showed that while effective at retrieving correct matches, it also retrieved a large number of search results (mostly false positives), increasing the likelihood of incorrect matches (Mohanty et al., 2020).
To better understand how these models differ in the context of historical photo identification, we conducted a small-scale benchmarking (see details in Appendix  B). This revealed notable differences between the two models: the new model consistently outperformed the old model in terms of average precision (i.e., ranking correct matches higher) and recall (i.e., finding more correct matches). While the new model retrieved significantly fewer results overall (an average of 20 compared to roughly 378 with the older model), those results were more accurate, with higher confidence scores for correct matches. The old model, in contrast, produced a large number of results but with more false positives, making it relatively less precise in this domain.
These differences in key metrics—such as the number of results, the ranking of more similar-looking candidates, and improved handling of diversity—make the old model and new model compelling candidates for this study. However, while these measurable differences exist, an important question remains: can users actually perceive these differences in practice? Since participants were unaware of which model was in use during each trial, the study explores whether users can rely on visible cues—such as fewer but more accurate results and better-ranked matches—to distinguish between the two models.
3.4.1. Latency in retrieving results
Prior work has demonstrated that latency—the time taken by a system to retrieve result—can significantly shape user perceptions of algorithmic systems, where even slight delays reduce engagement and satisfaction in search and chatbot contexts, but in some cases, slower responses may foster reflection or higher perceived quality (Park et al., 2019; Teevan et al., 2013; Brutlag, 2009; Gnewuch et al., 2018). In this study, we introduced an artificial delay to explore whether users consider latency when trying to distinguish between model changes and how it may influence their perception of model accuracy. Participants were assigned to one of four conditions: (i) Old Model retrieves results faster than New Model, (ii) New Model retrieves results faster than Old Model, (iii) both models retrieve results with random latency, or (iv) both models retrieve results with identical latency. For slower conditions, we introduced a 12-second delay, while standard conditions took 3 to 5 seconds.
3.5. Dataset
Our dataset consists of 20 diverse Civil War portraits sourced from the CWPS database, similar to prior studies using this dataset (Mohanty et al., 2020, 2019a). The pool includes images of Union and Confederate soldiers of different ranks, predominantly white males, as well as African American soldiers. We also added photos of two white females, an Asian, and a Hispanic soldier. Some photos have known positive matches in the database, while others do not, allowing us to capture varying model performance across a range of inputs. This selection reflects the over-representation of white males in the CWPS dataset, which stems from historical biases (Kusuma et al., 2022). Additionally, facial recognition models are known to have racial and gender biases, often performing less accurately on non-white and female faces (Buolamwini and Gebru, 2018; Raji and Buolamwini, 2019). As a result, we anticipated that these biases might influence participants’ perceptions of the models’ performance, especially in cases involving women and people of color.
For each participant, we randomly selected 7 photos from the 20-photo pool. These 7 photos were distributed across 8 trials, with one photo intentionally repeated using two different models in separate trials. This design facilitated a direct comparison between models for the same query photo, allowing participants to assess potential differences in model performance with a familiar visual reference. The search results, or potential candidates presented to participants in each trial, were retrieved by the facial recognition models from the CWPS database, ensuring that all photos came from the same search pool of relevant candidates.
3.6. Participants
We recruited a total of 252 participants globally (58.7% men and 41.3% women; mean age = 29.8 years, SD = 9.7 years) through Prolific (Prolific, 2024). Participants were all at least 18 years of age, fluent in English, and hailed from multiple countries. The countries with the most participants were South Africa (57 participants), Poland (46 participants), and the United Kingdom (39 participants). All participants completed an IRB-approved consent form before beginning the task. The study was approved by the university’s IRB.
3.7. Measurement
We collected several interaction and behavioral data points from 252 participants, each completing 8 trials, which resulted in a total of 2016 search sessions (see Table 1). In total, participants made 10646 comparison decisions across trials and completed 1764 pairwise model comparisons (7 per user). The new model was (randomly) assigned to 1015 search sessions, while the old model was assigned to 1001 sessions.
Measurement Description
Time spent on search tasks Overall time spent on the search results page and individual comparison times.
Comparison decisions Decisions made during side-by-side comparisons, with participants choosing between: facial match, replica, different people, or not sure.
Perceived accuracy of the model Slider ratings from -100 (very inaccurate) to 100 (very accurate) after each trial and an optional note justifying their decision.
Model comparisons Participants rated how similar or different the models were between two consecutive trials using a four-point Likert scale: very different, somewhat different, somewhat similar, or very similar.
Final model preference At the conclusion of the study, participants selected their preferred model between Model A-UP and Model B-UP for future tasks and provided a justification.
Table 1. Table summarizing the key measurements collected during the study, including participant behavior, decisions, and model evaluations over the 8 trials.
In order to understand the signals participants used to distinguish between models, we also calculated the following ”difference” metrics between consecutive trials for each participant, as participants compared these factors side-by-side during model comparisons (see Figure 4):
  • Difference in perceived accuracy: The change in participants’ perceived accuracy ratings (from -100 to 100) between two consecutive trials.
  • Difference in the number of search results retrieved: The difference in the number of results returned by the models in each trial, a key observable factor during model comparisons.
  • Difference in response latency: The time difference in the models’ response time (latency), as varying response times were introduced to explore their effect on participant perceptions.
To evaluate participants’ identification performance with the different models, we calculated the following metrics based on comparisons with the ground truth:
  • Precision: The proportion of correct identifications (true positives) out of all results flagged as a positive match by participants (i.e., classified as either a facial match or a replica). This measures how accurately participants identified correct matches.
  • Recall: The proportion of correct identifications (true positives) out of all actual positive matches available in the search results. This measures how well participants were able to find all the correct matches in the results retrieved by the model.
  • False Positive Rate (FPR): For trials where no correct matches were present in the results, FPR was calculated as the proportion of incorrect matches (false positives) classified by participants as either a facial match or a replica. This helps us assess how often participants mistakenly identified incorrect results as positive matches.
3.8. Analysis
To answer our research questions, we employed the following statistical methods:
  • Paired t-tests: Used to compare precision, recall, and false positive rates (FPR) between the models to assess improvements in identification performance. We also applied t-tests to examine participants’ ability to detect model changes and to compare perceived accuracy between the models.
  • Logistic mixed-effects models (GLMM): Used to predict participants’ ability to detect a model change (a binary outcome) and to analyze their final model preference. Key predictors included differences in perceived accuracy, latency, and search results between consecutive trials. GLMM was also applied to predict comparison decisions (e.g., positive vs. negative vs. uncertain decisions), using factors such as the underlying model type and time taken for individual comparisons. We implemented GLMM using the $glmer$ function in R (Bates et al., 2014).
  • Linear mixed-effects models (LME): Used to predict participants’ perceived accuracy of the models (a continuous outcome) and various behavioral metrics, including time taken for making comparison decisions, overall time spent browsing search results, number of results checked, and total number of comparisons. Key predictors included participants’ interactions with the models, such as the number of search results, the types of comparison decisions made (e.g., facial match or replica), and the underlying model type. We implemented LME using the $lmer$ function in R (Bates et al., 2014).
Mixed-effects models (both logistic and linear) were used to account for the nested structure of the data (multiple trials per participant), with random intercepts included to control for individual variability.
3.9. Findings
3.9.1. Participants struggled to distinguish the underlying facial recognition model used between two consecutive trials.
Out of 1764 pairwise model comparisons across 252 participants, the overall accuracy in detecting whether models had changed was 48.87%, close to random guessing. When the model changed, participants identified the change 56.98% of the time, but when the model remained the same, they accurately detected this only 38.38%. This difference in detection rates was statistically significant, with participants being better at identifying model changes than noticing when the model remained the same ($t(1665.5)=-7.91$, $p<0.001$). These findings contradict H1, as participants generally struggled to distinguish changes between models, especially when no change occurred.
To gain further insights, we examined participant comments after each trial. Some participants misinterpreted changes in the model’s behavior as variability or adaptations within a single model, rather than recognizing distinct models. One participant remarked: “Same AI? They got the 2 same matches,” after seeing similar outputs (e.g., common matches) across different trials, indicating that they thought they were interacting with the same model despite differences in other results. Another participant noted: “This model seems to be matching photos with consideration for ethnic features,” indicating they perceived the shift in behavior as an inherent adaptation of the same model, rather than a change to a different system.
3.9.2. Participants primarily relied on their perceived accuracy of the models rather than observable characteristics like latency or result count.
When comparing models across successive trials, participants were shown the latency, number of search results, and their own perceived accuracy for both trials. Differences in perceived accuracy played a key role in participants’ ability to distinguish between models. Specifically, for every one-unit increase in perceived accuracy difference, the likelihood of identifying the models as different increased substantially ($\beta=-0.89630$, $z=-12.122$, $p<0.001$). This indicates that as participants sensed greater differences in accuracy, they were much more likely to detect model changes.
On the other hand, differences in observable characteristics like latency and result count did not significantly affect participants’ ability to distinguish between models (latency: $\beta=0.06054$, $z=1.005$, $p=0.315$; result count: $\beta=0.03753$, $z=0.635$, $p=0.525$). These findings directly contradict H2, as participants primarily relied on their subjective assessment of model performance rather than observable characteristics such as latency or the number of search results when distinguishing models between trials.
3.9.3. Participants perceived the new model as more accurate than the old model, despite not knowing which model they were interacting with.
Across 2016 trials, participants rated the perceived accuracy of the models on a scale (see Figure 3) from -100 (highly inaccurate) to 100 (highly accurate). Participants were tasked with determining whether the model they used in each trial was the same as or different from the previous trial, but they were never informed whether they were interacting with the old model or the new model. As shown in Finding 3.9.1, participants struggled to distinguish whether the model changed between trials, meaning any potential conscious mapping to the actual underlying model (old or new) was effectively lost.
Despite this, participants consistently rated the new model as significantly more accurate (mean accuracy = 33.62) compared to the old model (mean accuracy = 18.32; $t(1981.5)=-6.10$, $p<0.001$), thus supporting H3. It is worth noting that neither model was perceived as highly accurate, with both models receiving moderate accuracy ratings on the scale (see Figure 6).
Figure 6. Participants perceived the new model to be more accurate compared to the old model despite moderate absolute accuracy ratings. The mean is denoted by the red dots in the boxplots.
[…]
Boxplot comparing perceived accuracy between two models: old model and new model. The new model’s boxplot shows higher perceived accuracy, with the mean indicated by a red dot. The new model also has a few outliers below -80, but overall participants rated it significantly higher than the old model.
3.9.4. User perceptions of AI model accuracy were largely shaped by their interactions with model outputs and engagement behaviors.
Figure 7. Factors influencing perceived accuracy. The plot shows effect size estimates (in units of perceived accuracy) for various factors. Positive factors such as facial match responses ($\beta=23.54$, meaning that each additional facial match response increases perceived accuracy by 23.54 points) and replica responses ($\beta=16.68$) had the largest impact on increasing perceived accuracy. In contrast, different person responses decreased accuracy ($\beta=-7.08$, meaning each additional response of this type reduces perceived accuracy by 7.08 points). Scrolling through more search results ($\beta=-4.75$) and time spent on the search page ($\beta=-6.29$) negatively impacted perceptions, while the number of search results retrieved had a small positive effect ($\beta=5.97$). Stars next to effect sizes denote statistical significance, with *** indicating $p<0.001$, ** indicating $p<0.01$, and * indicating $p<0.05$.
[…]
A dot-and-whisker plot showing the effect size estimates for various factors influencing perceived accuracy. Positive effects (in blue) include facial match responses ($\beta=23.54$) and replica responses ($\beta=16.68$), while negative effects (in red) include different person responses ($\beta=-7.08$) and scrolling through more search results ($\beta=-4.75$). Latency shows no significant effect ($\beta=0.38$). The chart has stars next to the effect size to indicate statistical significance.
We observed that when participants made positive comparisons, it influenced perceived accuracy in a positive way, with more facial match responses ($\beta=23.54,p<0.001$) and replica responses ($\beta=16.68,p<0.001$) leading to higher perceived accuracy (see Figure 7). In contrast, negative comparisons, such as identifying a Different Person, significantly reduced perceived accuracy ($\beta=-7.08,p<0.001$).
Participant comments reflected these findings, highlighting that personal experiences while comparing results strongly shaped their perceptions. One participant remarked, ”It appeared to retrieve a match as the top result but some of the other results looked nothing like the original person”, reinforcing how accuracy perceptions were influenced by direct comparison outcomes. Another noted, ”Facial hair seems to be a hindrance for facial recognition. Although some seemed slightly similar, none were correct”, pointing to specific features impacting perceived accuracy.
We also found that user behavior played an important role in shaping perceptions. Spending more time on individual comparisons positively influenced perceived accuracy ($\beta=9.64,p<0.001$). However, scrolling through more search results and spending additional time on the overall search page negatively affected perceived accuracy ($\beta=-4.75,p<0.001$; $\beta=-6.29,p<0.01$, respectively). Participants’ comments also pointed to concerns about diversity and representation in the model’s outputs, which influenced their perception of its accuracy. One participant stated, ”The selection of data it was trained on lacked validity because model output had more males than females in comparison to the female query photo”. Another noted, ”The race of the person on the photo and the race of the most similar matches the model suggested were different”, pointing to racial mismatches, which diminished their confidence in the model.
Finally, when participants interacted with the new model, they were more likely to rate the model as having higher perceived accuracy ($\beta=14.23,p<0.001$), even though they were unaware of which underlying model they were using. This is in line with H3. Interestingly, latency did not have a significant impact on perceived accuracy ($p=0.73$), while a higher number of search results had a small but significant positive effect on perceived accuracy ($\beta=5.97,p<0.001$).
3.9.5. User preferences were significantly driven by perceived model accuracy
Our analysis showed that perceived accuracy was a significant predictor of model preference. For every unit increase in the perceived accuracy of Model A-UP, the odds of preferring Model A-UP decreased by 3.4% ($\beta=-0.03398$, $z=-4.94$, $p<0.001$). Conversely, for every unit increase in the perceived accuracy of Model B-UP, the odds of preferring Model B-UP increased by 3.8% ($\beta=0.03799$, $z=5.69$, $p<0.001$). Other factors, such as the number of trials assigned to each model or whether a match was found, did not significantly influence model preference (all $p>0.05$).
3.9.6. Human-AI team performance showed minimal improvement in recall with the newer model, while precision and false positive rates remained similar.
When comparing participants’ decisions to ground truth across all 2016 search sessions, we observed only a marginal improvement in recall with the new model (62.64%) compared to the old model (56.05%), as confirmed by a two-sample Welch t-test ($t(1274.1)=-2.99$, $p=0.0028$). However, there was no significant difference in precision or false positive rates between the models (see Table 2). These findings suggest that LABEL:H4 was not fully supported, as improvements in Human-AI team performance with the more accurate model were minimal.
Models
At least 1 positive match
in the search results
No positive match
in the search results
Precision Recall False Positive Rate
Old Model 80.59% 56.05% 25.40%
New Model 81.10% 62.64% 27.5%
Table 2. User performance with the old and new facial recognition model across 2016 search sessions. The metrics used are mean precision, recall, and false positive rates.
3.9.7. The underlying model significantly influenced user behavior in terms of time spent and the kind of decisions made
While interacting with the outputs of the newer model, participants made 1 fewer comparison on average ($\beta=-1.00$, $t=-2.86$, $p<0.01$), spent 16.7 seconds less per session ($\beta=-16.74$, $t=-5.03$, $p=5.46\times 10^{-7}$), and scrolled through 11.6 fewer results ($\beta=-11.60$, $t=-13.38$, $p<2\times 10^{-16}$) compared to the older model.
Our findings further revealed that the underlying model significantly impacted both the decisions participants made and the time they spent engaging with the system. When interacting with the outputs of the new model, participants were 46% more likely to make positive decisions (i.e., classifying an image as a match or replica) compared to the old model ($\beta=0.38$, $z=7.76$, $p<0.001$), and 30.7% less likely to make negative decisions ($\beta=-0.37$, $z=-8.11$, $p<0.001$). The new model did not significantly affect uncertain decisions ($\beta=0.08$, $z=1.43$, $p=0.15$).
Participants also took 0.61 seconds longer per comparison with the new model ($\beta=0.61$, $t=3.72$, $p=2\times 10^{-4}$). Interestingly, the time spent per comparison was linked to decision type: longer times were associated with an increased likelihood of positive ($\beta=0.046$, $z=15.24$, $p<2\times 10^{-16}$) and uncertain decisions ($\beta=0.040$, $z=13.30$, $p<2\times 10^{-16}$), while shorter times were linked to negative decisions ($\beta=-0.099$, $z=-24.62$, $p<2\times 10^{-16}$). These findings support  H5, as the underlying model significantly influenced user behavior.
3.9.8. Perceived model changes had negligible impact on user behavior
Contrary to  H6, perceived model changes had a limited effect on user behavior. Participants who believed the model had switched did not significantly alter the number of comparisons made ($\beta=0.24$, $t=0.57$, $p=0.57$), the time spent per session ($\beta=-0.06$, $t=-0.015$, $p=0.99$), or the mean time spent per comparison ($\beta=-0.27$, $t=-1.04$, $p=0.30$). However, participants did scroll through 2 more results on average when they perceived a model switch ($\beta=2.04$, $t=1.98$, $p=0.048$), suggesting slight exploratory behavior.
3.10. Key Takeaways
Our findings revealed that despite significant performance differences between the two underlying models, participants were unable to accurately distinguish between them, contradicting H1. Although the newer model demonstrated higher precision and recall, participants did not achieve any performance gains, refuting H4. This lack of improvement suggests that participants did not adapt their strategies when interacting with the more accurate model. Had they been explicitly informed of the model changes, participants might have adjusted their behavior by focusing on fewer but more accurate results. For instance, they could have prioritized evaluating top-ranked candidates rather than exhaustively scrolling through large result sets, which would have aligned their approach with the strengths of the new model. Such an adjustment could have led to better task performance and more efficient decision-making.
Moreover, our analysis shows that participants relied heavily on their personal interactions with the system to form perceptions of accuracy, which then influenced how they distinguished between models. Rather than using objective indicators like latency or result count, users trusted their subjective experiences, highlighting the need to communicate system updates in ways that reflect how changes will affect user interactions, rather than simply presenting technical improvements. The failure of H6, where perceived model changes did not significantly alter user behavior, further underscores that participants may not understand how to adjust their strategies when models change. Raising awareness about model capabilities and providing explicit guidance on how to adapt could help users optimize their performance and make better use of updated system functionalities.
Interestingly, while we observed similar false positive rates and human-AI team performance across both models, the newer model did induce different user behaviors. Participants spent more time deliberating when making positive matches or uncertain decisions and were quicker to rule out negative matches. While this did not result in overall performance gains, it suggests that the new model may be encouraging more thoughtful decision-making. Despite minimal improvements in measurable outcomes, the new model appears to be pushing users toward more ”reflective” behaviors, which could be beneficial with further guidance or in longer-term use. Overall, this underscores the need for explicitly informing users about model updates, which could lead to more effective use of the new model’s capabilities and a clearer understanding of how these changes influence both behavior and perceptions.
4. Study 2: Comparing User Perceptions of Old and New AI Models in a Real-World Deployment
Study 1 revealed that when model updates were invisible, users struggled to detect changes and did not fully capitalize on the more accurate model’s potential, leading to only marginal improvements in human-AI team performance. In this study, we explore whether explicitly informing users of a new model would lead to more effective use of its outputs and how their preferences and perceptions would differ when directly comparing two models.
We investigated how users of Civil War Photo Sleuth (CWPS)—a platform where the older facial recognition model had been in use—perceived and compared a newly introduced AI model. These users, already familiar with the older model, were informed that a ”new model” was available, but were given no specific details about its characteristics or improvements. Unlike Study 1, where model updates were hidden from participants, this study allowed users to explicitly toggle between the old and new models in a real-world environment, giving them the agency to explore differences between the models on their own terms.
Our goal was to uncover not only user preferences but also the factors influencing these preferences and the folk theories users developed about the models’ performance in a practical, real-world setting.
4.1. Study Setup
(a) Results from the old model: 598 results retrieved.
[…]
The search interface of Civil War Photo Sleuth showing results from the old model. The interface displays 598 retrieved results with images of historical figures, including Theodore S. Bowers, Seager S. Atwell, Powell F. Clayton, and an unidentified person.
(b) Results from the new model: 28 results retrieved.
[…]
The search interface of Civil War Photo Sleuth showing results from the new model. The interface displays 28 retrieved results with images of historical figures, including Seager S. Atwell, Frederick A. Lull, Thomas J. Jackson, and Warren K. Snow.
Figure 8. Comparison of search results retrieved by the old model (a) and the new model (b). The old model retrieved a significantly larger number of results (598) compared to the new model (28). In addition to fewer results, the new model presents different people, suggesting improvements in the relevance of the results retrieved.
In April 2023, the CWPS team integrated the most recent facial recognition model (i.e., the new model from Study 1) alongside the existing older model (see Section 3.4 for model details). Users could seamlessly switch between the outputs of both models via a toggle button added to the search interface, allowing them to continue their regular photo investigation tasks (see Figure 8).
We recruited 10 CWPS users who had actively used the platform’s facial recognition-based search workflow within the last six months. The group comprised nine males and one female, with an average age of 46 (min = 19, max = 62). Each participant expressed a certain degree of familiarity with the platform’s facial recognition search feature. Participants investigated the identities of 10 photos of their choice, either by uploading new images or selecting from existing ones on the site.
Over two weeks, they were instructed to use both models and document their reflections in a diary (see template in Appendix), including which model they found more useful and why, along with a usefulness rating from -100 (not at all useful) to 100 (highly useful). A mid-point reflection was collected after five photo investigations, followed by a final reflection at the conclusion of the study to assess their overall preferences and perceptions of both models. This study was approved by the university’s IRB.
4.2. Analysis
We reviewed and categorized participants’ diary entries, initially identifying 28 distinct codes, which were later grouped into 7 main themes (”Preference”, ”Perceived Accuracy”, ”Result Counts”, ”Other performance factors”, ”Misc. visual factors”, ”Unrelated to model”, ”Behavior”) discussed in the findings section.
4.3. Findings
4.3.1. The new model was generally preferred, but not unanimously.
While most participants favored the new model, citing its accuracy and relevance in matching images (7 out of 10 participants), the older model remained useful in certain cases. As shown in Figure 9(a), participants chose the new model for 54 photos and the old model for 27. In the words of P10: ”The New Model is a clear winner in my book. I tried to challenge it using images that I knew contained the same person but with different angles and conditions and it almost always passed. The potential for matching faces in outdoor group images to studio portraits is tremendous here, and I think this will prove to be an invaluable tool.”
(a) Model Preferences of Users.
[…]
The stacked horizontal bar chart illustrates the preferences of 10 participants (P1 to P10) regarding the use of old and new AI models, as well as options for using neither or both in tandem. Each bar represents a participant and is divided into sections representing the number of times they chose each option. The data indicates a diverse range of preferences across participants. Participants P4, P5, and P9 showed a strong preference for the new model, with P10 exclusively preferring it. P8 notably leaned towards the old model, while P1 demonstrated a balanced preference with a slight lean towards the old model. Participants P2, P3, P6, and P7 exhibited mixed preferences, with some even opting for using both models or neither at various points. Overall, the chart portrays a nuanced picture of user preferences, indicating that different individuals have varied approaches and levels of comfort with the old and new AI models.
(b) Usefulness Scores for Old and New Models.
[…]
The bar chart compares the mean usefulness scores between the old and new models as rated by each participant (P1 to P10). The chart displays two sets of bars for each participant: one representing the old model (in sky blue) and the other representing the new model (in salmon). Each bar signifies the mean score, while the error bars indicate the standard deviation, showcasing the variability in scores given by each participant. The participants are listed along the x-axis, and the mean scores, ranging from -100 to 100, are plotted along the y-axis, with grid lines at intervals of 25 units.
Figure 9. (a) Model Preferences of Users and (b) Usefulness Scores for Old and New Models in the Diary Study.
Some users did not notice substantial differences between the models. As P1 observed, ”…I noticed little difference between the two. Both were fast, both produced similar results, as in the same possible people.” This was reflected in comparable scores for usefulness, as seen in Figure 9(b). Moreover, there were 19 instances where participants opted for neither model exclusively, instead choosing both. P7 suggested, ”Using both models in tandem is the way to go rather than comparing one directly to the other.” This sentiment was echoed by P2, who preferred toggling between models to explore different perspectives.
4.3.2. Cognitive load from sifting through large numbers of search results shaped user preferences.
Participants had mixed reactions to the number of search results, which significantly influenced their model preferences. Some users valued the larger result set from the old model for offering more choices and the potential for serendipitous discoveries. As P1 noted, ”There were more of a revolving door of options [with the old model], whereas the new model only gave me five.” P8 echoed this, appreciating the exploratory aspect: ”The fact that the image could possibly be buried way down the results page in the older model makes searching much more enjoyable.”
However, many participants found the larger result set overwhelming, citing the increased cognitive load when sifting through numerous results. P3 explained, ”After scrolling through the first 50, you kinda give up on looking for any more close matches.” P6 preferred the new model for its streamlined results, stating, ”It produces much fewer results, so I am not scrolling through an infinite number of possible matches with the old model.” Similarly, P10 found the old model overwhelming, remarking, ”The Old Model yielded 574 results (vs. just 17 from the New), so it was a bit overwhelming.”
Overall, participants who favored efficiency and lower cognitive load gravitated towards the new model, while those who valued more options and exploratory searches preferred the old model.
4.3.3. Perceived accuracy shaped user preferences differently across models
Participants’ preferences were often shaped by their perceptions of each model’s accuracy, though this varied on a case-by-case basis. For instance, P3 favored the new model for its accuracy in capturing facial features, noting, ”Hairlines, eyes, ears, facial all seem to match better.” On the other hand, P8 preferred the older model for its broader image characteristics, observing, ”Seems that with the old model, the resulting images had characteristics closer to my image: chin structure, hairstyle, eye colors.”
Some participants observed that both models were adept at finding matches, though their performance varied slightly in terms of accuracy. P5 shared an instance where both models successfully identified an alternate view of the same soldier, stating, ”Both models were able to find an alternate view of the same soldier. Even though this one is not inscribed, I can say without a doubt that it’s the same soldier as another image of Andrew Johnson. Results were similar in that the different view of Johnson showed up as the top result in both the new and old model.” However, P5 gave a slight edge to the new model, noting that it showed ”more facial similarity.” This sentiment was also observed by P6 and P9 in similar cases.
Participants also shared instances where one model clearly outperformed the other. P4 observed that the old model found a match the new model missed: ”That face turned up as #2 in the Old Model, whereas it didn’t even show up in the New Model.” Conversely, P6 shared a case where the new model succeeded in identifying a known subject that the old model missed: ”The new model correctly identified him with the very first subject.”
Ranking accuracy also played a role in how participants rated the models. P4 noted that, while both models retrieved a similar result, the new model ranked the match higher, stating, ”The New Model ranked the best matched photo highest, whereas the Old Model had it further down the queue.” In contrast, P5 found the old model to be more accurate in certain cases.
4.3.4. The trade-off between quantity and quality influenced user preferences for each model
Participants often experienced a trade-off between the number of search results and their perceived accuracy, which shaped their satisfaction with each model. The new model was frequently praised for its precision, as it provided fewer but more relevant results. P6 remarked, ”New model gave me much fewer matches again but those matches were much closer in appearance.” Similarly, P9 favored the new model, stating, ”I’d prefer the newer one because I am the type of person that prefers quality over quantity.”
However, the smaller pool of results also made inaccuracies more noticeable. P8 highlighted this issue, stating, ”The newer model always seemed to have at least 1/4 of the results that were not similar to the test image.” This led some users to prefer the old model for offering a larger pool of choices, despite its occasional inaccuracies. As P8 explained, ”The old model is way more preferable… it’s quite a bit more accurate and provides a larger results list.”
In contrast, participants like P3 were frustrated with the irrelevant matches that often came with the larger result set from the old model. P3 remarked, ”350 potential matches is too many to scroll through,” highlighting the tension between having more options and the cognitive load it created.
Ultimately, participants who valued precision favored the new model for its focused results, while those who preferred more choices gravitated towards the old model, even at the cost of sifting through irrelevant options.
4.3.5. Users proposed diverse and inconsistent folk theories about how both models work.
Category Old Model New Model
Facial Features “Stronger focus on the features of the face” (P7) “Wide range of characteristics…” (P8) “Results had all types of noses, including skinny ones that did not make sense” (P3) “Better job of finding facial matches with eyes, ears, mouths, etc.” (P5) “Focused on facial hair” (P9) “Focused on eyes and nose shape” (P10) “Focused on hairstyle first (hairstyles can change)” (P3)
Occlusion “Low resolution and large facial hair obstructing the face… hampers” (P10) “Obscured by shadow… seemed confused” (P10) “Seemed to focus on the Kepis they wore” (P3) “Puts more emphasis on headwear” (P5) “Low resolution and large facial hair obstructing the face… does not hamper as much” (P10)
Latency Perceptions “Felt model was slower” (P3) “Slower than the new model” (P8) “No difference in speed” (P9) “Model is faster as it produces fewer results” (P3) “No difference in speed” (P9)
Age Sensitivity “Showed more faces that didn’t look anything like the subject photo; for example, I was shown a photo of a very old man when the subject is obviously young to middle-aged” (P4) “I was amazed to see an older face matching the same face as a young man, as I’ve never seen that happen with the Old Model” (P10) “The new model gave me more results of youthful looking soldiers” (P5)
Non-facial Attributes “The model worked well with all image types that I tested (tintype, ambrotype, CDV)” (P8) “Picks up tagged words on the photos, not just the facial characteristics of the sitters” (P4)
Table 3. A comparison of user-formed folk theories regarding the old and the new facial recognition models. Quotes from participants illustrate divergent opinions on how each model interprets various aspects such as facial features, occlusions, latency perceptions, age sensitivity, and non-facial attributes.
Participants developed various folk theories to explain how the two models processed different aspects of the images. Their theories diverged significantly, often reflecting personal interpretations of the models’ behavior (see Table 3).
Facial Feature Recognition
Participants had differing views on how the models handled facial features. Some, like P5 and P9, believed the newer model excelled in matching specific facial details such as eyes and ears. P5 noted, ”I feel like the new model does a better job of finding facial matches with eyes, ears, mouths, etc.” Others, such as P8, felt the older model provided a broader analysis, stating, ”I could count on the old model producing results that incorporated a wide range of characteristics (beard style, mustache style, eyes, chin, cheeks, forehead, hairstyle, etc..).”
Occlusion and Headgear
Theories about occlusion and headgear varied across participants. P10 believed the newer model handled occlusions, such as shadows or facial hair, better than the older model: ”Images that are low resolution, poorly lit, or have large facial hair obstructing the face can all hamper this model somewhat, but not nearly as much as the Old Model.” Conversely, P5 and P3 speculated that the older model placed more emphasis on headgear, with P3 noting, ”This model seemed to focus on the Kepis that they wore.”
Latency Perceptions
Although both models had identical retrieval times, participants perceived the older model as slower due to the larger number of results it returned. P3 remarked, ”I felt this model was slower as it produced far more results.” In contrast, they found the newer model faster, adding, ”This model is faster as it produces fewer results.” P6, however, observed, ”Both models took about 2 seconds to correctly identify this subject.”
Age Sensitivity
Participants speculated that the models varied in their ability to match faces across different age groups. P4 observed that the older model struggled with age discrepancies, stating, ”The old model showed more faces that didn’t look anything like the subject photo; for example, I was shown a photo of a very old man when the subject is obviously young to middle aged.” In contrast, P10 praised the newer model for its ability to match faces across different ages: ”I was amazed to see an older face matching the same face as a young man, as I’ve never seen that happen with the Old Model.”
Non-Facial Attributes
Some participants believed the models incorporated non-facial attributes like image format or text inscriptions. P8 noted that the older model worked well across various image formats, stating, ”The model worked well with all image types that I tested (tintype, ambrotype, CDV).” P4 speculated that the newer model integrated Optical Character Recognition (OCR), observing, ”So now I see that the New Model picks up tagged words on the photos, not just the facial characteristics of the sitters.”
5. Discussion
5.1. Contextual Complexity in End-User Evaluation of AI Models
Despite the new AI model being more accurate, we did not observe unanimous support for it from the study participants. While some users clearly favored the newer model for its precision and fewer false positives, others preferred the older model due to its broader range of choices. This divergence highlights the complexity of user preferences when interacting with evolving AI models. Although there may be a possibility of neophilia (i.e., a bias towards new things or innovations) (Wikipedia contributors, 2022), our findings suggest that users were evaluating the models based on context-specific needs. Similar to other algorithm-driven platforms (Register et al., 2023; Milton et al., 2023; DeVito et al., 2018), participants crafted multiple, often divergent folk theories (see Table 3). These theories reflected attempts to understand the system’s inner workings, with users assessing the models based on how well they fit their specific use cases rather than reacting purely to novelty.
Users considered a wide array of factors in their evaluations, many of which shape the overall user experience and extend beyond traditional performance benchmarks. Factors such as cognitive load, search result volume, and the time required to sift through results were repeatedly mentioned, highlighting the importance of conducting exploratory evaluations that incorporate the actual user base. Standard technical metrics may not fully capture user experience, especially given significant individual differences. For example, participants who valued precision tended to favor the new model for its focused, high-accuracy results, whereas those who preferred a broader set of options gravitated toward the older model. To better capture these individual preferences, future research could benefit from larger-scale experiments tracking behavioral metrics, such as scrolling patterns and time spent evaluating results, and correlating these with user actions and preferences.
In both studies, participants quickly identified model shortcomings, but their evaluations were nuanced. In the first study, where users were not informed which model was in use, they struggled to differentiate between them. Even in the diary study, where model changes were explicit, inconsistent folk theories emerged, with no clear preference for one model. This contrast suggests that end-user evaluation of AI models is complex, influenced by individual needs and task context. Future work could explore long-term behavioral patterns to determine whether users continue toggling between models or eventually settle on a preferred one.
5.2. Deploying AI Model Updates in a Human-AI Collaborative Environment
Our findings underscored the importance of effectively communicating AI model updates in human-AI collaborative environments, especially when users interact with the outputs of the models on a regular basis. Across both studies, we saw that users place significant emphasis on their personal experiences, developing diverse and sometimes inconsistent folk theories about model behavior. These folk theories often diverge from standard benchmarking metrics, suggesting that conventional performance metrics alone may be insufficient to convey a model’s capabilities.
A more user-centered approach to deploying updates is necessary. One effective strategy, as evidenced in our second study, is to provide a ”try-it-on” or ”beta testing” interface that enables users to directly compare different models and explore their strengths and weaknesses. This approach allows users to identify changes firsthand and form more informed judgments. However, given the inconsistencies in folk theories, developers could also integrate social components that facilitate knowledge sharing. Enabling users to share insights and discuss their experiences could foster a more comprehensive understanding of model behavior and updates, leveraging the collective experiences of the user community, drawing social transparency ideas in AI systems (Ehsan et al., 2021).
A novel direction could involve developing platform-specific LLM agents capable of generating logs and reports of user interactions with the AI models. Aggregating these reports over time and across users could yield valuable insights into how models perform in different contexts, ultimately improving the deployment and adaptation process for both users and developers.
Finally, providing educational interfaces to help users understand key distinctions between different types of models (Kusuma et al., 2022)—such as deterministic versus stochastic behavior, which emerged in Study 1—can be crucial. This could also help users, who may not be AI-literate, to figure out key strategies on how to interact with different AI models. By combining exploratory interfaces, social learning opportunities, dynamic feedback mechanisms, and targeted education, developers can create a more supportive environment for users to effectively engage with and understand AI model updates.
6. Conclusion
In this work, we investigated how users distinguish and perceive different black-box AI models, specifically focusing on facial recognition models in the context of historical person identification. We first conducted a benchmarking study and found that the newer, developer-certified model performs better than the older model on a historical dataset. However, despite the discernible differences in capabilities, this did not translate to a significant performance gain in human-AI team performance with the newer model.
In the online experiment we conducted, we observed that crowd workers found it challenging to discern between the models across trials. Notably, they largely relied on their perceived accuracy of the models, instead of focusing on visible model characteristics such as the number of search results and latency. When real-world users of the Civil War Photo Sleuth (CWPS), a facial recognition-based platform for identifying historical photos, were asked to compare the two models, we received mixed responses. While the majority leaned towards the newer model, a subset preferred the older model, and a few expressed a desire to use both models in tandem. These preferences were influenced by various factors including perceived accuracy and cognitive load. Interestingly, users also developed several divergent folk theories surrounding each model’s capabilities.
As part of future work, we aim on exploring how individual differences and design interventions for communicating model differences can potentially influence the way people perceive and interact with different models. Overall, our work opens doors for future research in the area of AI-infused systems, offering insights into the potential for smoother model updates and fostering a more harmonious human-AI collaboration.
Appendix A Generative AI Usage
We used ChatGPT to 1) generate the initial captions and descriptions for the images, 2) for polishing the quality of text, and 3) formatting the tables.
Appendix B Benchmarking the Old and New Facial Recognition Models
From Table 4, we find that new model consistently outperforms old model across all evaluated metrics — average precision, recall, and reciprocal rank — on our study dataset. There are several instances where new model was able to retrieve correct matches at higher ranks when old model was unsuccessful. This indicates that new model has a more accurate and refined search capability compared to old model. In addition, new model tends to return a significantly smaller number of results compared to old model, which is indicative of the model’s precision.
We also observe that new model consistently retrieves correct matches at higher ranks and with greater confidence scores than old model (see Table 5). This confirms that new model not only identifies matches more accurately but also assigns them higher confidence scores, reflecting its enhanced performance. Importantly, new model also performs better across a diverse range of faces, including those of African American (Photo IDs: 42234, 25008), Hispanic (Photo ID: 41622), and female (Photo IDs: 41631, 41825) subjects.
Photo ID Average Precision Reciprocal Rank Recall Total Search Results
Old Model New Model Old Model New Model Old Model New Model Old Model New Model
52778 0.00 0.00 0.00 0.00 0.00 0.00 625 142
40914 0.54 1.00 1.00 1.00 0.80 0.80 785 35
51831 0.00 1.00 0.00 1.00 0.00 1.00 775 16
29626 0.35 1.00 1.00 1.00 0.75 0.75 789 17
41787 0.63 1.00 1.00 1.00 1.00 1.00 792 17
30357 0.00 1.00 0.00 1.00 0.00 1.00 763 12
34096 1.00 1.00 1.00 1.00 1.00 1.00 25 11
19561 1.00 1.00 1.00 1.00 1.00 1.00 846 15
46204 1.00 1.00 1.00 1.00 1.00 1.00 14 11
19595 0.00 0.00 0.00 0.00 0.00 0.00 144 15
29405 0.02 0.75 0.04 1.00 0.71 0.71 849 32
46459 0.00 1.00 0.00 1.00 0.00 1.00 24 13
42234 0.05 1.00 0.05 1.00 1.00 1.00 108 10
24322 1.00 1.00 1.00 1.00 1.00 1.00 26 5
25008 0.75 1.00 1.00 1.00 1.00 1.00 63 6
24356 0.00 0.00 0.00 0.00 0.00 0.00 38 11
41901 0.00 0.00 0.00 0.00 0.00 0.00 37 9
41622 0.00 1.00 0.00 1.00 0.00 1.00 695 11
41631 1.00 1.00 1.00 1.00 0.50 1.00 4 7
41825 0.92 1.00 1.00 1.00 1.00 1.00 156 7
Mean 0.41 0.79 0.50 0.80 0.54 0.76 377.90 20.10
Table 4. Model Benchmarking Results: Average precision, reciprocal rank, recall, and the total number of search results for all photos in the study dataset.
Photo ID Ranks of Correct Matches Confidence Scores of Correct Matches
Old Model New Model Old Model New Model
52778 [] [] [] []
40914 [1, 2, 33, 68] [1, 2, 3, 4] [80.11%, 78.98%, 68.78%, 67.32%] [87.85%, 84.79%, 77.08%, 76.39%]
51831 [] [1] [] [87.66%]
29626 [1, 92, 187] [1, 2, 3] [73.86%, 59.54%, 57.98%] [77.74%, 75.88%, 75.53%]
41787 [1, 2, 6, 107] [1, 2, 3, 4] [74.34%, 72.85%, 66.06%, 58.58%] [89.26%, 81.85%, 79.86%, 79.86%]
30357 [] [1] [] [74.61%]
34096 [1] [1] [63.01%] [91.24%]
19561 [1] [1] [77.25%] [90.22%]
46204 [1] [1] [59.25%] [91.27%]
19595 [] [] [] []
29405 [23, 87, 149, 179, 791] [1, 2, 3, 8, 20] [61.34%, 58.43%, 56.72%, 56.09%, 51.08%] [81.44%, 79.48%, 71.06%, 64.86%, 52.23%]
46459 [] [1] [] [75.50%]
42234 [19] [1] [54.78%] [83.86%]
24322 [1] [1] [78.66%] [93.84%]
25008 [1, 4] [1, 2] [70.51%, 59.59%] [89.62%, 76.02%]
24356 [] [] [] []
41901 [] [] [] []
41622 [] [1] [] [75.02%]
41631 [1, 2, 3] [1, 2, 3, 4, 5, 6] [69.19%, 68.91%, 59.14%] [92.18%, 92.11%, 90.06%, 78.26%, 78.01%, 75.02%]
41825 [1, 2, 4] [1, 2, 3] [70.10%, 64.84%, 61.56%] [80.13%, 79.60%, 75.09%]
Table 5. Model Benchmarking: Ranks and confidence scores of correct matches retrieved by Model 1 and Model 4.
Appendix C Mixed Effects Model Outputs
C.1. Distinguishing Models: Perceived Accuracy vs. Objective Characteristics
We include the detailed output of the generalized linear mixed effects model used to examine whether participants relied on observable model characteristics (e.g., latency, result count) or their own perceived accuracy to distinguish between models. This analysis corresponds to Section 3.9.2 in the main text.
Generalized linear mixed model fit by maximum likelihood (Laplace Approximation) [’glmerMod’]
Family: binomial ( logit )
Formula: user_model_comparison_binary ~ latency_diff_scaled + result_count_diff_scaled + perceived_accuracy_diff_scaled + (1 | session_id)
Data: df
AIC BIC logLik deviance df.resid
2159.2 2186.5 -1074.6 2149.2 1752
Scaled residuals:
Min 1Q Median 3Q Max
-1.7272 -0.7875 -0.3763 0.8605 6.2023
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 0.4657 0.6825
Number of obs: 1757, groups: session_id, 252
Fixed effects:
Estimate Std. Error z value Pr(>|z|)
(Intercept) -0.47974 0.07086 -6.770 1.29e-11 ***
latency_diff_scaled 0.06054 0.06022 1.005 0.315
result_count_diff_scaled 0.03753 0.05908 0.635 0.525
perceived_accuracy_diff_scaled -0.89630 0.07394 -12.122 < 2e-16 ***
[…]
Correlation of Fixed Effects:
(Intr) ltnc__ rsl___
ltncy_dff_s -0.005
rslt_cnt_d_ 0.003 -0.321
prcvd_ccr__ 0.222 -0.027 0.059
C.2. Factors influencing Perceived Accuracy
We include the detailed output of the generalized linear mixed effects model used to examine which factors influenced the perceived accuracy of a given model. This analysis corresponds to Section 3.9.4 in the main text.
Linear mixed model fit by REML. t-tests use Satterthwaites method [’lmerModLmerTest’]
Formula: perceived_accuracy ~ facial_match_count + rep_match_count + ns_count + diff_count + search_results_count + model_type + overall_timer +
seen_tiles + mean_target_comparison_time + latency + (1 | session_id) + (1 | task_id)
Data: df
REML criterion at convergence: 20852.3
Scaled residuals:
Min 1Q Median 3Q Max
-5.2073 -0.4940 0.1214 0.6061 5.3309
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 555.61 23.571
task_id (Intercept) 24.96 4.996
Residual 1554.75 39.430
Number of obs: 2016, groups: session_id, 252; task_id, 8
Fixed effects:
Estimate Std. Error df t value Pr(>|t|)
(Intercept) 19.1127 2.7291 21.9115 7.003 5.10e-07 ***
facial_match_count 23.5355 1.0502 1998.9006 22.411 < 2e-16 ***
rep_match_count 16.6758 1.0244 1984.1011 16.279 < 2e-16 ***
ns_count 2.9388 1.0867 2000.0761 2.704 0.006902 **
diff_count -7.0787 1.5766 1906.5875 -4.490 7.55e-06 ***
search_results_count 5.9723 1.2112 1884.9401 4.931 8.91e-07 ***
model_typerecognition_04 14.2293 2.3095 1842.6505 6.161 8.84e-10 ***
overall_timer -6.2852 2.0136 1505.9057 -3.121 0.001834 **
seen_tiles -4.7527 1.2930 1976.4139 -3.676 0.000244 ***
mean_target_comparison_time 9.6421 1.4457 1931.4743 6.669 3.34e-11 ***
latency 0.3769 1.0954 2000.2490 0.344 0.730837
[…]
Correlation of Fixed Effects:
(Intr) fcl_m_ rp_mt_ ns_cnt dff_cn srch__ md__04 ovrll_ sn_tls mn_t__
fcl_mtch_cn 0.051
rp_mtch_cnt 0.035 0.074
ns_count 0.036 0.056 0.106
diff_count 0.015 0.121 -0.042 0.006
srch_rslts_ -0.223 -0.070 -0.072 -0.099 0.004
mdl_typr_04 -0.426 -0.119 -0.083 -0.084 -0.036 0.524
overall_tmr -0.007 -0.183 -0.144 -0.262 -0.647 0.042 0.016
seen_tiles -0.073 0.025 0.040 0.010 -0.139 -0.097 0.172 -0.335
mn_trgt_cm_ 0.019 0.119 0.096 0.087 0.414 -0.072 -0.044 -0.552 0.195
latency -0.002 0.010 0.024 0.045 0.110 -0.313 0.006 -0.162 0.008 0.070
C.3. Factors influencing model preference
We include the detailed output of the logistic regression model used to analyze the factors influencing participants’ preferences between Model A-UP and Model B-UP. The analysis evaluates the effects of trial count, match findings, and perceived accuracy on model preference, discussed in Section 3.9.5. The significant predictors are highlighted in the output below.
glm(formula = preferred_model_binary ~ model_A_trial_count +
model_A_trial_count_match_found + model_A_avg_perceived_accuracy +
model_B_trial_count_match_found + model_B_avg_perceived_accuracy,
family = binomial(link = "logit"), data = df)
Coefficients:
Estimate Std. Error z value Pr(>|z|)
(Intercept) 0.400568 1.327662 0.302 0.763
model_A_trial_count -0.308623 0.299720 -1.030 0.303
model_A_trial_count_match_found 0.075469 0.189530 0.398 0.690
model_A_avg_perceived_accuracy -0.033977 0.006874 -4.942 7.71e-07 ***
model_B_trial_count_match_found 0.092030 0.190740 0.482 0.629
model_B_avg_perceived_accuracy 0.037990 0.006681 5.686 1.30e-08 ***
[…]
(Dispersion parameter for binomial family taken to be 1)
Null deviance: 321.43 on 240 degrees of freedom
Residual deviance: 225.78 on 235 degrees of freedom
(11 observations deleted due to missingness)
AIC: 237.78
Number of Fisher Scoring iterations: 5
C.4. Impact of the underlying model on the comparison decisions
We include the detailed output of the generalized linear mixed effects model used to examine the impact of the underlying model on the comparison decisions. This analysis corresponds to Section 3.9.7 in the main text.
These are the outputs for the three models: positive vs. rest, negative vs. rest, and uncertain vs. rest.
Generalized linear mixed model fit by maximum likelihood (Laplace Approximation) [’glmerMod’]
Family: binomial ( logit )
Formula: positive_vs_rest ~ model_type + (1 | session_id)
Data: df
AIC BIC logLik deviance df.resid
11472.1 11493.9 -5733.0 11466.1 10643
Scaled residuals:
Min 1Q Median 3Q Max
-2.1231 -0.6113 -0.4218 0.7908 5.4912
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 0.8151 0.9028
Number of obs: 10646, groups: session_id, 252
Fixed effects:
Estimate Std. Error z value Pr(>|z|)
(Intercept) -1.01287 0.06717 -15.080 < 2e-16 ***
model_typerecognition_04 0.38091 0.04906 7.764 8.21e-15 ***
[…]
Correlation of Fixed Effects:
mdl_typr_04 -0.378
Generalized linear mixed model fit by maximum likelihood (Laplace Approximation) [’glmerMod’]
Family: binomial ( logit )
Formula: negative_vs_rest ~ model_type + (1 | session_id)
Data: df
AIC BIC logLik deviance df.resid
13100.2 13122.0 -6547.1 13094.2 10643
Scaled residuals:
Min 1Q Median 3Q Max
-3.6717 -0.8580 0.4317 0.7729 3.0907
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 0.8162 0.9034
Number of obs: 10646, groups: session_id, 252
Fixed effects:
Estimate Std. Error z value Pr(>|z|)
(Intercept) 0.18270 0.06562 2.784 0.00537 **
model_typerecognition_04 -0.36777 0.04533 -8.113 4.92e-16 ***
[…]
Correlation of Fixed Effects:
mdl_typr_04 -0.338
Generalized linear mixed model fit by maximum likelihood (Laplace Approximation) [’glmerMod’]
Family: binomial ( logit )
Formula: uncertain_vs_rest ~ model_type + (1 | session_id)
Data: df
AIC BIC logLik deviance df.resid
8750.3 8772.1 -4372.1 8744.3 10643
Scaled residuals:
Min 1Q Median 3Q Max
-0.9543 -0.4570 -0.3581 -0.2578 4.9061
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 0.5016 0.7082
Number of obs: 10646, groups: session_id, 252
Fixed effects:
Estimate Std. Error z value Pr(>|z|)
(Intercept) -1.81609 0.06201 -29.287 <2e-16 ***
model_typerecognition_04 0.08289 0.05790 1.432 0.152
[…]
Correlation of Fixed Effects:
mdl_typr_04 -0.465
C.5. Impact of Model Type on Search Behavior
We include the detailed output of the linear mixed effects models used to examine the impact of the underlying model on the search behavior, specifically on the number of comparisons made, the time spent on the search results page, and the number of results checked. This analysis corresponds to Section 3.9.7 in the main text.
Linear mixed model fit by REML. t-tests use Satterthwaites method [’lmerModLmerTest’]
Formula: total_num_comparisons ~ model_type + (1 | session_id)
Data: df
REML criterion at convergence: 14139.9
Scaled residuals:
Min 1Q Median 3Q Max
-2.0686 -0.1943 -0.0892 -0.0029 20.8738
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 8.64 2.939
Residual 59.12 7.689
Number of obs: 2016, groups: session_id, 252
Fixed effects:
Estimate Std. Error df t value Pr(>|t|)
(Intercept) 5.7866 0.3079 518.8372 18.791 < 2e-16 ***
model_typerecognition_04 -1.0048 0.3509 1896.8430 -2.863 0.00424 **
[…]
Correlation of Fixed Effects:
mdl_typr_04 -0.574
Linear mixed model fit by REML. t-tests use Satterthwaites method [’lmerModLmerTest’]
Formula: overall_timer ~ model_type + (1 | session_id)
Data: df
REML criterion at convergence: 23280.3
Scaled residuals:
Min 1Q Median 3Q Max
-2.7329 -0.4519 -0.1463 0.2863 18.4644
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 1510 38.86
Residual 5248 72.44
Number of obs: 2016, groups: session_id, 252
Fixed effects:
Estimate Std. Error df t value Pr(>|t|)
(Intercept) 125.079 3.378 427.941 37.032 < 2e-16 ***
model_typerecognition_04 -16.744 3.331 1856.454 -5.027 5.46e-07 ***
[…]
Correlation of Fixed Effects:
mdl_typr_04 -0.497
Linear mixed model fit by REML. t-tests use Satterthwaites method [’lmerModLmerTest’]
Formula: seen_tiles ~ model_type + (1 | session_id)
Data: df
REML criterion at convergence: 17784.4
Scaled residuals:
Min 1Q Median 3Q Max
-1.9846 -0.4864 -0.0937 0.1384 13.1841
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 53.69 7.327
Residual 360.65 18.991
Number of obs: 2016, groups: session_id, 252
Fixed effects:
Estimate Std. Error df t value Pr(>|t|)
(Intercept) 21.0858 0.7632 516.8290 27.63 <2e-16 ***
model_typerecognition_04 -11.5951 0.8669 1895.9167 -13.38 <2e-16 ***
[…]
Correlation of Fixed Effects:
mdl_typr_04 -0.572